OpenStudy (kaylacat4):

The amount that results when $3,000 is compounded at 7% annually over eight years. ?? 1792.44 4680 5154.56

OpenStudy (mathmate):

The compound interest formula: Future value = P(1+i)^n i=annual interest, compounded annually, expressed in decimal n=number of years. P=principal Example: 500 at 3% interest compounded annually for 10 years: FV=500(1.03)^10=671.96

OpenStudy (kaylacat4):

thank you!